OpenStudy (anonymous):

Which is the equation of the line that contains points (8, 10) and (–4, 2)?

OpenStudy (anonymous):

use slope formula (y2-y1)/(x2-x1)

OpenStudy (anonymous):

\[x-x1/x2-x1=y-y1/y2-y1\]

OpenStudy (anonymous):

here x1=8 y1=10 and x2=-4 and y2=2

OpenStudy (anonymous):

after finding the slope use y-y1=(slope)*(x-x1)